OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the graph of the rational function. Show your work. f(x) = x-4/x(x-4)

OpenStudy (anonymous):

x=1 is your vertical asymptote though if there was a factor of x-1 on top then it could possibly be a hole that is f(x)=(x-1)/(x-1) has a hole at x=1 (not a vertical asymptote) that is f(x)=x(x-1)/(x-1) also has a hole at x=1 but g(x)=(x-1)/(x-1)^2 has a vertical asympote at x=1 since it has more (x-1)'s on bottom then on top my answer Teacers reply>You find the VA's by setting the denominator to 0. When we get a factor that cancels out, it is not a vertical asymptote. What is it? Graph it to see what it looks like

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Is this from a test or something?

OpenStudy (anonymous):

no its from my work

OpenStudy (anonymous):

i havent taken my midterm yet lol

OpenStudy (jack1):

f(x) = x-4/x(x-4) is that: \(\Large f(x) = x- \frac4x \times(x-4) \) or \(\Large f(x) = x- \frac4{x \times(x-4)} \) or \(\Large f(x) = \frac{x-4}{x \times(x-4)} \)

OpenStudy (anonymous):